What is the speed of an electron with kinetic energy 730 eV ?

What is the speed of an electron with kinetic energy 0.38 keV ?

Thank you!
Answer #1

A.)

Given Kinetic energy(KE) = 730 eV = 730*(1.6*10^-19) J

KE = 1.168*10^-16 J

also kinetic energy is given by,

KE = 0.5*m*v^2

here, m = mass of electron = 9.1*10^-31 kg

v = velocity of electron = ??

So,

0.5*(9.1*10^-31)*v^2 = 1.168*10^-16

v^2 = (2*1.168*10^-16)/(9.1*10^-31)

v = 1.60*10^7 m/s

B.)

Given Kinetic energy(KE) = 0.38 keV = 0.38*10^3 eV = (0.38*10^3)*(1.6*10^-19) J

KE = 6.08*10^-17 J

also kinetic energy is given by,

KE = 0.5*m*v^2

here, m = mass of electron = 9.1*10^-31 kg

v = velocity of electron = ??

So,

0.5*(9.1*10^-31)*v^2 = 6.08*10^-17

v^2 = (2*6.08*10^-17)/(9.1*10^-31)

v = 1.16*10^7 m/s